# ONEBench to Test Them All: Sample-Level Benchmarking Over Open-Ended Capabilities

### v1

@@ Abstract
Traditional fixed test datasets fall short in evaluating the open-ended capabilities of foundation models. To address this limitation, we propose ONEBench (OpeN-Ended Benchmarking), a new testing paradigm that consolidates individual evaluation datasets into a unified, ever-expanding sample pool. ONEBench allows users to generate custom, open-ended evaluation benchmarks from this sample pool, corresponding to specific capabilities of interest. By aggregating and reusing samples across test sets, ONEBench enables the assessment of diverse capabilities beyond those covered by the original test sets, while mitigating overfitting and dataset bias. Most importantly, it frames model evaluation as a collective process of selecting and aggregating sample-level tests.
[…]
@@ 1 Introduction
ONEBench to rule them all (evals),
ONEBench to find them (samples),
ONEBench to test them all (models),
and by queried ability rank them.
@@
Deep learning has arrived in the post-dataset era [^1] From a talk by Alexei Efros at ICML 2020. With the rapidly expanding range of zero-shot capabilities of foundation models, the focus of model evaluation has moved beyond singular, dataset-specific performance measurements obtained by splitting a fixed collection of data into training and test sets. Instead, foundation models are employed as general knowledge and reasoning engines across all domains in which they prove to be useful. This creates a pressing need to characterize their open-ended set of capabilities across various metrics in zero-shot settings [39], for flexible evaluation over several use-cases. However, traditional static benchmarks, which test generalization on fixed test splits, are unable to effectively probe the ever-evolving set of capabilities of foundation models. This raises an important question: How can benchmarking adapt to measure an open-ended set of capabilities?
We propose a solution based on dynamic sample-level evaluation, which we call ONEBench for OpeN-Ended Benchmarking, where test sets for particular capabilities are generated ad-hoc from a large pool of individual annotated data samples. These sample-level evaluations act as atomic units of measurement that can be combined into an exponential variety of aggregations. Due to this flexibility, the sample pool and corresponding annotation metrics can be continuously updated to include new evaluations. Additionally, this approach can reduce dataset bias—systematic quirks in the data arising from how it was collected [151, 87, 171]. Finally, by combining samples across test sets, ONEBench can better capture real-world diversity [111].
The most important feature of ONEBench is the potential to democratize evaluation. Unlike traditional benchmarks, typically created by individual groups based on their own criteria for data collection and evaluation procedures [27, 7], ONEBench allow the integration of test sets from multiple sources reflecting a wide range of perspectives, use-cases, and objectives. This flexibility allows different interest groups with varying needs to collaboratively define their own evaluations by selecting the most appropriate combination of tests to suit their specific requirements. Moreover, the design of ONEBench challenges the dominant approach of chasing single benchmark scores, which do not indicate the difficulty of individual data instances [32], in favour of a plurality of rankings and dynamic, granular, multi-faceted evaluation.
Challenges in ONEBench. To build ONEBench, we must address two main challenges: (a) Heterogeneity and (b) Incompleteness. Heterogeneity means that model measurements span different metric types, such as binary (correct/incorrect), numeric (BLEU scores), and ordinal (preference rankings). This diversity makes it difficult to aggregate measurements and standardize comparisons across different models. Incompleteness, on the other hand, arises from models being evaluated on different, unequal subsets of testing data, rendering direct aggregation unfair and inaccurate. Traditional benchmarks use a multi-task setting, where each component test set requires evaluating all models over a fixed set of samples using a single metric, completely sidestepping both these issues.
@@
Figure 1: The ONEBench Framework. Left: ONEBench comprises a set of models, a pool of data samples spanning multiple test sets, metadata describing models and data samples, and a collection of heterogeneous, sample-level measurements. Right: the user formulates a query to capture the desired model capability, using a mix of structured metadata filters and semantic search. Selected models are then ranked on a subset of data samples that meet the specified criteria.
@@
Solution and Theoretical Guarantees. To tackle these challenges, we apply social choice theory, treating data samples as voters who express preferences among models. By converting all measurements into ordinal rankings, we leverage well-established principles to create a robust algorithm for aggregating over heterogeneous and incomplete data. Our approach assumes a random utility model based on the Plackett-Luce framework, which provides guarantees for accurately recovering ground-truth utility scores from input measurements. This approach ensures that our model rankings are both theoretically sound and practical, with rapid convergence guarantees enabling accurate rankings from relatively small amounts of data.
@@
Empirical Validation. We develop two instantiations of ONEBench: ONEBench-LLM for language models and ONEBench-LMM for multimodal(vision and language) models. These benchmarks unify evaluations across their respective domains by aggregating data from diverse sources, from arena-style human preference data [18, 91] to heterogeneous multi-task leaderboards [8, 79, 176, 23]. Our empirical results demonstrate that the Plackett-Luce model [119, 93] is a good fit for aggregating real-world benchmarks, showing a high correlation with ground-truth score-based rankings over homogeneous datasets. Importantly, we demonstrate that this strong correlation holds even when up to $95$% of the data is missing. This robustness allows us to reduce costs by 2020{20}20 times with little loss in performance by selecting a subset of data samples for evaluation. Finally, we compare Plackett-Luce rankings with widely adopted ranking metrics like ELO [30] and Bradley-Terry [15] and show that our method outperforms them on accuracy and robustness to missing information.
@@
Personalized Aggregation. Consider the following scenario: you are a scientist in a biochemistry lab and require an LLM to assist with designing experiments related to antibodies. With ONEBench, you can input a query, such as “immunology” or “antibodies”, and receive a dynamically constructed benchmark that ranks models based on their performance on this specific capability. While the optimal selection of personalized capability sets is an emerging research field, we present a proof of concept by distinguishing between tasks (e.g., reading comprehension) and concepts (e.g., Clostridium bacteria). We show how a combination of structured filters and flexible semantic search allows users to define their capability of interest with respect to these two dimensions and perform targeted evaluations resulting in personalized model rankings.
@@
In essence, ONEBench represents a democratized, open-source collection of diverse evaluation samples enriched with detailed metadata, driven by an aggregation method designed to robustly rank models across heterogeneous metrics and incomplete evaluation data. Users can conduct semantic searches and apply structured query filters to dynamically generate a benchmark tailored to their specific use case. They can also contribute new evaluation samples and sample-level model measurements, which can then be instantly aggregated, producing personalized rankings. This framework enables lifelong aggregation of arbitrary test sets with unprecedented flexibility and precision.
@@ 2 ONEBench: Formulation
At the heart of ONEBench is the idea of homogenizing performance evaluation across benchmarks by replacing benchmark-specific metrics with rankings. Importantly, this can be done at the level of individual data samples. In the following, we describe the process of construction and evaluation in detail, together with mathematical guarantees.
@@ 2.1 Components
The goal of building ONEBench (${\mathcal{B}}$) from a growing set of benchmarks $({\mathcal{B}}_{k})_{k=1}^{N}$ is to evaluate a collection of models (${\mathcal{M}}$) using an ever-expanding test pool of data instances ($\mathcal{D}$) which may be annotated with additional meta-data specifying the capabilities ($\mathcal{C}$) tested. To cope with the diversity of data originating from different benchmarks, sample-level rankings (${\mathcal{S}}$) are created for all data instances in the test pool. We provide a schematic overview of ONEBench in Fig. 1 and describe each component below:
@@
i) Pool of Data. ${\mathcal{D}}{=}{((x_{1},y_{1}),\dots,(x_{n},y_{n}))}$ denotes an ordered collection of test data instances $x_{i}$ with annotation $y_{i}$. An example of a data instance $x_{k}$ is the question “What was the dominant strain of Flu in 2010? Select among the four choices.” with the reference answer “H1N1/09” represented by $y_{k}$. In addition, information about tested capabilities can be provided as metadata, for example as a list of keywords such as temporal Q&A, pandemics, history, biology, virology, multiple-choice Q&A, beyond the specific dataset it originates from. Typically, the data samples are obtained via pooling from $N$ different benchmarks $({\mathcal{B}}_{k})_{k=1}^{N}$ and we refer to the subset of data instances obtained from benchmark ${\mathcal{B}}_{k}$ as ${\mathcal{D}}_{k}\subseteq\mathcal{D}$.
ii) Models. ${\mathcal{M}}{=}{(f_{base},f_{1},\dots,f_{m})}$ is a set of $m+1$ models, whose capabilities are evaluated with respect to a baseline $f_{base}$. An example of $f_{base}$ is a random model. The original benchmarks $({\mathcal{B}}_{k})_{k=1}^{N}$ will likely cover different sets of models ${\mathcal{M}_{\mathcal{B}_{k}}}\subseteq\mathcal{M}$ for their evaluations.
iii) Sample-level Rankings. For each data instance $(x_{j},y_{j})\in\mathcal{D}$ a sample-level ranking ${s_{j}\in\mathcal{S}}$ is created for the subsets of models ${\mathcal{M}_{j}:=\mathcal{M}\cap\mathcal{M}_{\mathcal{B}_{k(j)}}}$ where $k(j)$ denotes the index of the benchmark from which the data instance $(x_{j},y_{j})$ was collected. Importantly, sample-level rankings are a function of the metrics used by the different benchmarks that discard any information about the specifics of the metrics. This is the key to our approach of enabling the aggregation across heterogeneous evaluation paradigms and metrics. More specifically, $s_{j}\in\mathcal{S}$ represents an ordinal ranking over the models $\mathcal{M}_{j}$ for sample $(x_{j},y_{j})$ represented by a permutation $\sigma_{j}$ such that $f_{\sigma_{j}(1)}\succeq\dots\succeq f_{\sigma_{j}(m_{j})}$ where $m_{j}=|\mathcal{M}_{j}|$ is the number of models compared in the $j$-th sample-level ranking. In addition, for each $k$ we distinguish the case $f_{\sigma(k-1)}\succ f_{\sigma(k)}$ if $f_{\sigma(k-1)}$ performs better than $f_{\sigma(k)}$ and $f_{\sigma(k-1)}\sim f_{\sigma(k)}$ in case of indistinguishable performance. Thus, each sample-level ranking $s_{j}\in\mathcal{S}$ can be uniquely determined by a mapping $\sigma_{j}:\{1,\dots,m_{j}\}\to\{1,\dots,m\}$ with $\sigma_{j}(k)$ providing the index of the model in $\mathcal{M}$ that is on the $k$-th place in the ordering for the $j$-th sample-level ranking and $\pi_{j}\in\{\succ,\sim\}^{m_{j}-1}$ defining the corresponding binary sequence of pairwise performance relations.
Ordinal Rankings and Information Loss. Using ordinal measurements leads to information loss, which can impede downstream aggregation algorithms due to the data processing inequality ([148], Section 2.8). This principle asserts that any estimation made from processed data cannot outperform estimation based on the original, unprocessed data. However, cardinal measurements frequently suffer from calibration issues, even within a single metric [134]. Consequently, in practice, ordinal measurements can paradoxically outperform cardinal ones despite the inherent information loss.
@@
iv) Capabilities. To support the selective retrieval of all relevant sample-level rankings in $\mathcal{B}$ based on user queries, it is possible to endow the sample-level rankings with additional capability $c\in\mathcal{C}$. Of course, modelling the range of capabilities that different evaluators might be interested in is a research challenge in itself. Here, we only provide a proof of concept, for which we define two categories of capabilities: (1) tasks, like multiple-choice question answering, captioning, and translation, and (2) concepts like makeup, geometry, tarantula, etc. The rationale for this broad interpretation in ONEBench is to explore which capabilities can be reliably evaluated in a dynamic, flexible manner through queries. Since the set of capabilities is open-ended, we only tag the data samples with task information. The relevant samples for a given concept are retrieved from ONEBench dynamically at test time through semantic search.
@@
Lifelong Expansion of ONEBench. The data pool ($\mathcal{D}$) and model IDs ($\mathcal{M}$) are stored as a table, while sample-level model measurement ($\mathcal{S}$) are represented as a relational database linking these two tables. Constructing a lifelong heterogenous benchmark augments $\mathcal{D}$, $\mathcal{M}$ and $\mathcal{S}$ with the following operations: ${\mathcal{B}}{=}{(\mathcal{D},\mathcal{M},\mathcal{S},\text{{insert}}_{\mathcal{D}},\text{{insert}}_{\mathcal{M}}},\text{{insert}}_{\mathcal{S}}$). Operations $\text{{insert}}_{\mathcal{D}}$ and $\text{{insert}}_{\mathcal{M}}$ for expanding the data pool are straightforward: add new samples and new models to the corresponding table. The $\text{{insert}}_{\mathcal{S}}$ operation inserts a sample-level ranking. Additional metadata is saved to enable retrieval over database rows sharing the same original metric, such as ‘BLEU score’, or ‘exact match’.
2.2 Querying Capabilities for Personalized Evaluation
@@
To evaluate a given capability, ONEBench takes a dynamic approach. First, we retrieve ($\texttt{retrieve}_{\mathcal{D}}$) samples that match the query. Then, we aggregate ($\texttt{aggregate}_{\mathcal{S},\mathcal{D}}$) the sample-level rankings to produce the overall ranking.
@@
Retrieve ($\texttt{retrieve}_{\mathcal{D}}$). In this step, the system selects relevant data instances based on a user’s query. The query language is flexible and allows retrieving data instances that semantically relate to a specific topic or match certain criteria. The retrieval is implemented through a combination of k-nearest neighbors (kNN) search on dense embeddings using the query as the input and structured queries that take advantage of the unified data schema. We provide extensive empirical analysis to test retrieval quality.
@@
Aggregate ($\texttt{Aggregate}_{\mathcal{S},\mathcal{D}}$). The sample-level measurements over the retrieved subset of data samples is combined using the random utility modelling approach [163], which defines a joint probability distribution over all measurements assuming statistical independence:
$p(s_{1},\dots,s_{n_{\infty}}|\gamma_{1},\dots,\gamma_{m})=\prod_{j=1}^{n_{\infty}}p(s_{j}=[.]_{(\sigma_{j},\pi_{j})}|\gamma_{1},\dots,\gamma_{m}).$
@@
$p\left(s_{j}=[.]_{(\sigma_{j},\pi_{j})}\right)=\frac{\gamma_{\sigma_{j}(1)}}{\underbrace{\sum_{k=1}^{m_{j}}\gamma_{\sigma_{j}(k)}}_{f_{\sigma_{j}(1)}}}\times\frac{\gamma_{\sigma_{j}(2)}}{\underbrace{\sum_{k=2}^{m_{j}}\gamma_{\sigma_{j}(k)}}_{f_{\sigma_{j}(2)}}}\times\cdots\times\frac{\gamma_{\sigma_{j}(m_{j}-1)}}{\underbrace{\gamma_{\sigma_{j}(m_{j}-1)}+\gamma_{\sigma_{j}(m_{j})}}_{f_{\sigma_{j}(m_{j})}}},$
defining one parameter $\gamma_{k}$ for each model $f_{k}$ that determines its performance relative to all other models. To aggregate the model performances over all sample-level rankings, we determine the parameters
$\hat{\gamma}_{1},\dots\hat{\gamma}_{m}=\operatorname*{argmax}_{(\gamma_{1},\dots\gamma_{m})\in\mathbb{R}^{m}}\log p(s_{1},\dots,s_{n_{\infty}}|\gamma_{1},\dots,\gamma_{m})$
with maximum likelihood estimation. The global ranking is given by the permutation $\sigma_{\infty}$ for which $\hat{\gamma}_{\sigma_{\infty}(1)}>\dots>\hat{\gamma}_{\sigma_{\infty}(m)}$. The maximum likelihood condition uniquely determines all performance parameters $\{\hat{\gamma}_{k}\}_{k=1}^{m}$, as the likelihood function is strictly concave. The parameters of the Plackett-Luce model are identifiable up to an arbitrary additive constant. Consistency and asymptotic normality can also be shown under certain assumptions about the comparison graph [50]. We refer to the estimated latent variables $\{\hat{\gamma}_{k}\}_{k=1}^{m}$ as model scores. A model with a higher score is more likely to perform better on a randomly picked sample-level task than one with a lower score. To fix the arbitrary additive constant, we set the score of the baseline model $\hat{\gamma}_{baseline}$ to zero.
3 Aggregation in ONEBench: Theory and Practice
We view aggregating sparse ordinal preferences over models through a computational social choice lens—samples are voters, models are candidates, and the aggregation algorithm is the voting mechanisms [16]. Using established methods, we aggregate ordinal comparisons with partial data to produce a global ranking and analyze the properties of this resultant ranking.

### v2

@@ Abstract
Traditional fixed test datasets fall short in evaluating the open-ended capabilities of foundation models. To address this, we propose ONEBench (OpeN-Ended Benchmarking), a new paradigm that consolidates individual evaluation datasets into a unified, ever-expanding sample pool. ONEBench enables custom benchmarks for specific capabilities while reusing and aggregating samples, mitigating overfitting and dataset bias for broader capability assessment. It reframes model evaluation as selecting and aggregating sample-level tests. Transitioning from task-specific benchmarks to ONEBench introduces two challenges: heterogeneity (aggregating diverse metrics) and incompleteness (comparing models tested on different data subsets). To address these, we propose an aggregation algorithm that ensures identifiability—asymptotically recovering ground-truth scores—and rapid convergence, enabling accurate model comparisons with relatively little data. On homogenous datasets, our algorithm produces rankings that highly correlate with average scores. Moreover, it remains robust to over $95$% missing measurements, reducing evaluation costs by up to 20 times. We introduce ONEBench-LLM for language models and ONEBench-LMM for vision-language models, enabling targeted model testing across diverse capabilities.
ONEBench to Test Them All:
Sample-Level Benchmarking Over Open-Ended Capabilities
@@ 1 Introduction
Deep learning has arrived in the post-dataset era [^1] From a talk by Alexei Efros at ICML 2020. As foundation models rapidly expand their zero-shot capabilities, the focus of evaluation has moved beyond singular, dataset-specific performance measurements that rely on dividing fixed collections of data into train and test sets. Instead, foundation models are employed as general knowledge and reasoning engines across a wide range of domains. This creates a pressing need to characterize their open-ended capabilities using diverse metrics in zero-shot settings (Ge et al., 2024). However, static benchmarks, which test generalization on fixed test splits, cannot probe the ever-evolving capabilities of foundation models effectively. This raises an important question: How can benchmarking adapt to measure an open-ended set of capabilities?
@@
Figure 1: The ONEBench Framework. Left: ONEBench comprises a set of models, a pool of data samples spanning multiple test sets, metadata describing models and data samples, and a collection of sample-level measurements. Right: the user formulates a query to capture the desired model capability, using a mix of structured metadata filters and semantic search. Selected models are then ranked on a subset of data samples that meet the specified criteria.
We propose a solution based on dynamic, sample-level evaluation, which we call ONEBench (OpeN-Ended Benchmarking). In this approach, test sets for particular capabilities are generated ad-hoc from a large pool of individual annotated data samples. These sample-level evaluations act as atomic units of measurement that can be flexibly aggregated into an exponential number of configurations. Thanks to this flexibility, the sample pool and corresponding annotation metrics can be continuously updated to incorporate new evaluations. Additionally, this approach can reduce dataset bias—systematic quirks in the data arising from its collection process (Liu and He, 2024). Finally, by combining samples across test sets, ONEBench captures real-world diversity (Ni et al., 2024).
The most important feature of ONEBench is its potential to democratize evaluation. Unlike traditional benchmarks, typically created by individual groups based on their own criteria for data collection and evaluation procedures Bansal and Maini (2024), ONEBench integrates test sets from multiple sources reflecting a wide range of perspectives, use cases, and objectives. This flexibility allows different interest groups to collaboratively define their own evaluations by selecting the most appropriate combination of tests that best suit their specific requirements. Moreover, the design of ONEBench challenges the dominant approach of chasing single benchmark scores, which fail to account for the difficulty of individual data instances (Ethayarajh et al., 2022), in favor of a plurality of rankings and a dynamic, granular, multi-faceted evaluation.
Challenges. Building ONEBench requires addressing two key challenges: heterogeneity and incompleteness. Heterogeneity arises because model evaluations span diverse metric types, such as binary (correct/incorrect), numeric (BLEU scores), and ordinal (preference rankings), making aggregation difficult. Incompleteness occurs when models are tested on non-overlapping subsets of data, preventing fair and direct comparisons. Traditional benchmarks sidestep these issues by using a multi-task setup, where all models are evaluated on the same samples using a single metric.
@@
Solution and Theoretical Guarantees. We address these challenges using social choice theory, treating data samples as voters expressing preferences over models. By converting all measurements into ordinal rankings, we leverage established principles to robustly aggregate heterogeneous and incomplete data. Our approach assumes a random utility model based on the Plackett-Luce framework (Plackett, 1975; Luce, 1959), which provides guarantees for accurately recovering ground-truth utility scores. This approach ensures that our model rankings are both theoretically sound and practical, with rapid convergence guarantees enabling accurate rankings from limited data.
Why rankings? While converting cardinal measurements to ordinal ones results in information loss, two key considerations justify this choice in benchmarking: external validity and benchmark longevity. External validity (Liao et al., 2021) refers to how well evaluation results generalize across different settings. Recht et al. (2019) found that although accuracy shifts across test sets, model rankings remain stable. Salaudeen and Hardt (2024) extended this by comparing ImageNet models to those trained and tested on a new dataset, arguing that ordinal comparisons are more robust, e.g., injecting random labels into two otherwise identical benchmarks causes accuracy to vary while preserving relative order. Given the noise in LLM and LMM evaluation (from prompt design, label choices, in-context examples, etc.), rankings offer a more reliable signal than absolute values. Benchmark longevity is another growing challenge, as internet-scale datasets lead to faster saturation. Contamination and adaptive overfitting threaten evaluation integrity. While ONEBench mitigates contamination by filtering compromised samples, its use of rankings helps guard against overfitting—an idea inspired by the Ladder algorithm (Hardt and Recht, 2022).
@@
Empirical Validation. We create ONEBench for two domains: ONEBench-LLM for language models and ONEBench-LMM for vision-language models. These benchmarks unify evaluations by aggregating data from diverse sources, including preference data (arenas) and heterogeneous multi-task leaderboards. Our empirical results demonstrate that the Plackett-Luce model effectively aggregates real-world benchmarks, showing a high correlation with ground-truth score-based rankings over homogeneous datasets. Notably, this strong correlation persists even when up to $95$% of the data is missing, enabling a 20 times reduction in evaluation costs with minimal impact on performance. Finally, we compare Plackett-Luce rankings to widely adopted methods such as ELO Elo (1967) and Bradley-Terry (Bradley and Terry, 1952), demonstrating superior accuracy and robustness to missing data. We additionally compare our Plackett-Luce rank aggregation approach to popular computational social choice theory methods like Borda Count and the Dowdall System (section 3.2.4).
@@
Personalized Aggregation. Imagine you are a biochemist seeking an LLM to assist with designing experiments related to antibodies. With ONEBench, you can input a query, such as “immunology” or “antibodies” to generate a dynamically constructed benchmark that ranks models based on their performance in this specific domain. While the optimal selection of personalized capability sets remains an open research challenge, we present a proof of concept by distinguishing between tasks (e.g., reading comprehension) and concepts (e.g., Clostridium bacteria). By combining structured filters and flexible semantic search, users can define their capability of interest along these dimensions and conduct targeted evaluations, resulting in personalized rankings.
@@
To summarize, ONEBench is a democratized, open-source collection of diverse evaluation samples enriched with detailed metadata. Its robust aggregation method ranks models across heterogeneous metrics and incomplete evaluation data. Users can perform semantic searches and apply structured query filters to dynamically generate benchmarks tailored to their needs. They can also contribute new evaluation samples and model measurements, which are instantly aggregated to refine rankings. This framework enables lifelong aggregation of arbitrary test sets with unprecedented flexibility and precision.
@@ 2.1 Components
The goal of ONEBench is to evaluate a set of models $\{m_{k}\}_{k=1}^{M}$ using a continuosly expanding pool of test data samples $\mathcal{D}$ drawn from multiple benchmarks $\{\mathcal{B}_{k}\}_{k=1}^{B}$. Each data sample may include metadata specifying the capabilities it is testing. To handle the diversity of data from different benchmarks, we generate sample-level rankings (${\mathcal{S}}$) for all samples in the test pool. Figure 1 provides a schematic overview of ONEBench, with each component described below.
@@
(i) Data Pool. The data pool $\mathcal{D}=\{(x_{k},y_{k})\}_{k=1}^{D}$ consists of data samples $x_{k}$ with reference answers $y_{k}$. An example of a data sample is the question “What was the dominant strain of flu in 2010? Select among four choices." with a reference answer “H1N1/09". Each instance can also include metadata specifying tested capabilities, for example as a list of keywords like temporal QA, pandemics, history, biology, virology, multiple-choice QA.
(ii) Models. The set of models is defined as =$\mathcal{M}=\{m_{base}\}\cap\{m_{k}\}_{k=1}^{M}$, where $m_{base}$ serves as a baseline for evaluating the capabilities of the other models. A common choice for $m_{base}$ is a random model. Since the original benchmarks evaluate different sets of models, each benchmark ${\mathcal{B}_{k}}$ considers a subset of models ${\mathcal{M}_{\mathcal{B}_{k}}}\subseteq\mathcal{M}$.
(iii) Sample-level Rankings. For each data sample $(x_{j},y_{j})\in\mathcal{D}$, we construct a sample-level ranking ${s_{j}\in\mathcal{S}}$ over model subset ${\mathcal{M}_{j}\subseteq\mathcal{M}_{\mathcal{B}_{k}}}$, where $k$ denotes the index of the benchmark from which the sample $(x_{j},y_{j})$ was collected (more in appendix C). Crucially, these rankings depend only on the evaluation metrics used by each benchmark, abstracting away the specifics of those metrics. This abstraction is central to our approach, as it enables aggregation across heterogeneous evaluation metrics.
@@
(iv) Capabilities. To enable selective retrieval of relevant sample-level rankings in $\mathcal{B}$ based on user queries, each ranking can be associated with a capability. We distinguish between tasks (e.g., question answering, captioning) and concepts (e.g., makeup, geometry). Since capabilities are inherently open-ended, we only tag data samples with task information, while concept-based retrieval is performed dynamically at test time using semantic search.
@@
Lifelong Expansion. Data pool $\mathcal{D}$ and model set $\mathcal{M}$ are stored as tables, while sample-level model evaluations are represented as a relational database linking these tables. Expanding ONEBench over time requires augmenting $\mathcal{D}$, $\mathcal{M}$, and $\mathcal{S}$ through the following operations: $\texttt{insert}_{\mathcal{D}},\texttt{insert}_{\mathcal{M}},\texttt{insert}_{\mathcal{S}}$. The first two operations simply add new samples and models, while $\texttt{insert}_{\mathcal{S}}$ registers a new sample-level ranking.
2.2 Capability Querying
@@
To evaluate a given capability, ONEBench takes a dynamic approach. First, we retrieve samples that match the query. Then, we aggregate the sample-level rankings to produce the overall ranking.
@@
Retrieve ($\texttt{retrieve}_{\mathcal{D}}$). Here, the system selects relevant data instances based on a user’s query. The query language is flexible and allows retrieving data instances that semantically relate to a specific topic or match certain criteria. Retrieval combines k-nearest neighbors (kNN) search over dense embeddings, with the query as input, and structured queries that leverage the unified data schema.
@@
Aggregate ($\texttt{Aggregate}_{\mathcal{S},\mathcal{D}}$). Measurements over the retrieved subset are combined using the random utility modelling approach (Xia, 2019), defining a joint probability distribution over all measurements (sample rankings $s_{j}$ and model scores $\gamma_{j})$, given model permutations $\sigma_{j}$ and binary sequence of pairwise performance relations $\pi_{j}$ assuming statistical independence:
$p(s_{1},\dots,s_{n_{\infty}}|\gamma_{1},\dots,\gamma_{M})=\\ \prod_{j=1}^{n_{\infty}}p(s_{j}=[.]_{(\sigma_{j},\pi_{j})}|\gamma_{1},\dots,\gamma_{M}).$ (1)
@@
$p\left(s_{j}=[.]_{(\sigma_{j},\pi_{j})}\right)=\\ \frac{\gamma_{\sigma_{j}(1)}}{\underbrace{\sum_{k=1}^{m_{j}}\gamma_{\sigma_{j}(k)}}_{f_{\sigma_{j}(1)}}}\times\cdots\times\frac{\gamma_{\sigma_{j}(m_{j}-1)}}{\underbrace{\gamma_{\sigma_{j}(m_{j}-1)}+\gamma_{\sigma_{j}(m_{j})}}_{f_{\sigma_{j}(m_{j})}}},$ (2)
defining one parameter $\gamma_{k}$ for each model $m_{k}$ that determines its performance relative to all other models. To aggregate model performances over sample rankings, we estimate parameters
$\hat{\mathbf{\gamma}}=\operatorname*{argmax}_{\mathbf{\gamma}\in\mathbb{R}^{m}}\log p(\mathbf{s}\mid\mathbf{\gamma})$
with maximum likelihood estimation (MLE). The global ranking follows the permutation $\sigma_{\infty}$ where $\hat{\gamma}_{\sigma_{\infty}(1)}>\dots>\hat{\gamma}_{\sigma_{\infty}(m)}$. The ML condition uniquely determines all performance parameters $\{\hat{\gamma}_{k}\}_{k=1}^{M}$, as the likelihood function is strictly concave. The parameters of the Plackett-Luce model are identifiable up to an arbitrary additive constant. Consistency and asymptotic normality can also be shown under certain assumptions about the comparison graph (Han and Xu, 2023). We refer to the estimated latent variables $\{\hat{\gamma}_{k}\}_{k=1}^{M}$ as model scores. A model with a higher score likely performs better on a randomly picked sample-level task than one with a lower score. To fix the additive constant, we set the baseline model score $\hat{\gamma}_{baseline}$ to zero.
3 ONEBench: Aggregation
We view aggregating sparse ordinal preferences over models through a computational social choice lens, where samples are voters, models are candidates, and the aggregation algorithm is the voting mechanism (Brandt et al., 2016). We aggregate ordinal comparisons with partial data to produce a global ranking and analyze its properties.